# Benchmarks and applications of the nuclear deexcitation event generator NucDeEx

### v2

@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant No. 23KJ0319. The author expresses deep gratitude to the host researcher of this JSPS fellowship, Professor Yoshinari Hayato, for his invaluable assistance in improving the manuscript. The author is grateful to the SAMURAI-79 Collaboration for valuable discussions on the deexcitation process and its simulations. The author also thanks to Dr. Yota Hino for the use of NucDeEx in the simulation.
@@ Appendix A Energy spectra of decay particles
Figs. 710 shows predicted energy spectra of decay particles from $^{11}\text{B}^{*}$, $^{11}\text{C}^{*}$, $^{15}\text{N}^{*}$, and $^{15}\text{O}^{*}$. The same excitation energy as Sec. III is assumed. Although it depends on the particles and generators, it is observed that 40–50% of decay particles have energies below the threshold. Furthermore, not only the absolute total emission probability, but also the shape of the spectra vary largely depending on the generators. The gamma-ray energy spectra predicted by INCL++/ABLA is above  8 MeV, indicating that low-lying discrete excitated states are not taken into account.
Figure 7: Energy spectra of decay particles from $^{11}\text{B}^{*}$ with excitation energy of 16–35 MeV predicted by deexcitation event generatoprs. The vertical dashed lines represent the detection energy thresholds in the experiment at RCNP [55, 56, 42]. The numbers shown in each pannel represent the emission probability of each particle: total emission probability and that above the threshold are given in this order for each event generator. All decays both of single-step and multi step are included. This emission probability can exceed 100% when multiple particles are emitted.

### v3

@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant No. 23KJ0319 and 25H00631. The author expresses deep gratitude to the host researcher of this JSPS fellowship, Professor Yoshinari Hayato, for his invaluable assistance in improving the manuscript. The author is grateful to the SAMURAI-79 Collaboration for valuable discussions on the deexcitation process and its simulations. The author also thanks to Dr. Yota Hino for the use of NucDeEx in the simulation.
@@ Appendix A ENERGY SPECTRA OF DECAY PARTICLES
Figures 710 shows predicted energy spectra of decay particles from $^{11}\text{B}^{*}$, $^{11}\text{C}^{*}$, $^{15}\text{N}^{*}$, and $^{15}\text{O}^{*}$. The same excitation energy as Sec. III is assumed. Although it depends on the particles and generators, it is observed that 40%–50% of decay particles have energies below the threshold. Furthermore, not only the absolute total emission probability, but also the shape of the spectra vary largely depending on the generators. The gamma-ray energy spectra predicted by INCL++/ABLA is above  8 MeV, indicating that low-lying discrete excited states are not taken into account.
Figure 7: Energy spectra of decay particles from $^{11}\text{B}^{*}$ with excitation energy of 16–35 MeV predicted by deexcitation event generators. The vertical dashed lines represent the detection energy thresholds in the experiment at RCNP [60, 61, 47]. The numbers shown in each panel represent the emission probability of each particle: total emission probability and that above the threshold are given in this order for each event generator. All decays both of single-step and multistep are included. This emission probability can exceed 100% when multiple particles are emitted.